# Visual Hallucinations of Multi-modal Large Language Models

### v1

@@ 1 Introduction
A multi-modal LLM (MLLM) [27, 30, 1, 6, 23] generates a text response for a given image and question. An MLLM typically comprises three components: a vision encoder, a vision-language connector, and an LLM. The vision encoder (e.g., CLIP [18]) converts an image into an embedding vector. The vision-language connector projects an image embedding vector into the LLM’s word embedding space. The projected vector is concatenated with the token embeddings of the question to form an input to the LLM, which generates a text response.
However, MLLMs often generate text responses containing factually incorrect details about an image, known as visual hallucination (VH) [10, 13]. Figure 1 shows an example where the MLLM hallucinates two lamps, contradicting the three lamps in the image. VHs in MLLMs pose obstacles to developing safe and trustworthy AI, which is emphasized in a recent U.S. Executive Order calling for rigorous testing to address potential harms from advanced AI systems [22].
@@
Prior works have tried to benchmark MLLMs’ VHs related to object existence [10, 12], optical character recognition (OCR), object counting, object positions comparing [3], orientation, and viewpoint [25] (concurrent to ours). However, they collect VH images only from existing image datasets like COCO [11]. This limits the diversity of VH images since they can only find a limited number of them. Moreover, existing image datasets may have been used to pre-train an MLLM, leading to data contamination [7, 21]. As a result, such VH images lead to a biased understanding of an MLLM’s performance, e.g., an MLLM is incorrectly concluded to perform well under VH.
@@
We propose VHTest, a tool that generates VH instances to test MLLMs. A VH instance is a triple (an image, a question, a reference answer). Our VHTest has three key steps. Step I finds initial VH instances using existing image datasets like COCO. Specifically, we first identify image pairs with high CLIP embedding similarity but low DINO v2 [16] embedding similarity. Such image pairs have contradictory similarities from two powerful vision encoders, indicating potential VHs. We note that Step I is also used in a concurrent work [25]. We then manually design questions and reference answers for these images to obtain initial VH instances. Step II generates a text description for a VH mode using the initial VH instances. A text description describes visual properties of a VH image. Finally, Step III uses a text-to-image generative model (e.g., DALL-E 3) to generate new images based on the text descriptions. Moreover, based on some templates, we design questions and reference answers for the generated images to construct VH instances.
Using our VHTest, we construct a new benchmark dataset for evaluating VHs in MLLMs. Our benchmark contains 1,200 VH instances covering 8 VH modes. The 8 VH modes are related to existence, shape, color, orientation, OCR, size, position, and counting of visual objects in an image. Note that shape and size VH modes are formulated by us, while the other 6 VH modes were also considered in prior [27] and concurrent [25] studies. Figure 2 and Figure 7 show some VH instances generated by VHTest.
@@ 3.1 Step I: Finding Initial VH Instances
Since the CLIP vision encoder is the backbone of many popular MLLMs such as LLaVA [14], LLaMA-Adapter [4], and mPLUG-Owl [28], we leverage CLIP to find initial VH instances. Our goal is to find images in an existing image dataset (e.g., COCO) that are incorrectly embedded by the CLIP vision encoder to have high similarity, despite differences in visual semantics. Such images may lead to VH for MLLMs due to their incorrect embeddings.
Specifically, given an image pair $(x_{1},x_{2})$, we compute the cosine similarity between their CLIP embedding vectors, i.e., $cos(f_{C}(x_{1}),f_{C}(x_{2}))$, where $f_{C}(\cdot)$ is an embedding vector output by CLIP. Moreover, we also use DINO v2 [16] as a reference vision encoder to compute their embedding vectors and compute their cosine similarity, i.e., $cos(f_{D}(x_{1}),f_{D}(x_{2}))$, where $f_{D}(\cdot)$ is an embedding vector output by DINO v2. Then, we find image pairs that have large cosine similarity under CLIP but small cosine similarity under DINO v2. In particular, we find image pairs that satisfy $cos(f_{C}(x_{1}),f_{C}(x_{2}))\geq 0.9$ and $cos(f_{D}(x_{1}),f_{D}(x_{2}))\leq 0.55$. Such image pairs are our candidates.
@@ 3.2 Step II: Text Description Generation
Summary version
PROVIDE: {image, question, hallucinated response, detected hallucination in the response} PROMPT:   [Through the example, explain the potential causes of such hallucination and how to produce more images and questions.]
@@
Summary version
PROVIDE: {image, question, hypothetical hallucinated response, detected hallucination in the hypothetical response} PROMPT:   [Through the example, explain the potential causes of the hallucination and how to generate more images and questions.]
@@ 3.3 Step III: Generating More VH Instances
Given a VH image, we prepare a question based on it. Specifically, we leverage object-driven templates to create diverse and relevant VH questions. For instance, “Describe the shape of the [object] in the picture.” is a template for the shape VH mode. We curate question templates for each VH mode and they are shown in Appendix C. Given a template for a VH mode, a human worker generates a question via manually analyzing the objects in the VH image. For instance, the human worker may replace the [object] as “pear” in the template above when the VH image contains a pear. The human worker verifies that the question should have a non-ambiguous answer based on the VH image. If no questions with non-ambiguous answers can be constructed, the VH image is discarded.
@@ 3.4 Benchmark Construction
We use VHTest to build two benchmarks. Specifically, we find initial VH instances in COCO [11] and we generate 150 VH instances for each VH mode, which results in 1,200 VH instances across 8 VH modes in total. This benchmark consists of “open-ended question” (OEQ), which requires manually labeling the responses of MLLMs when testing them on the benchmark. Therefore, to facilitate automatic evaluation when testing MLLMs, we also construct a closed-ended “yes/no question” (YNQ) version of the benchmark.
Specifically, for each VH instance, we convert the open-ended question into a binary “yes/no” question. For instance, the open-ended question “Describe the shape of the pear in the picture.” is converted into “Is the shape of the pear in the picture a square?”. Moreover, the reference answer is converted into yes or no. We construct the binary questions to ensure that the YNQ benchmark has a 50/50 split of “yes”/“no” reference answers.
@@ 4 Experiments
4.1 Experimental Setup
MLLMs
We evaluate three state-of-the-art MLLMs on our benchmarks: GPT-4V with its “gpt-4-vision-preview” version, LLaVA-1.5-13b [14], and MiniGPT-v2  [2].
Evaluation Metric
4.2 Testing VH in MLLMs
Initial VH Instances are Insufficient
Table 1 shows the accuracy of GPT-4V on the 100 initial VH instances from COCO for each VH mode, and the number of successful initial VH instances in each VH mode. We observe that a large fraction of the initial VH instances are not successful. In particular, the average accuracy of GPT-4V across the 8 VH modes is 0.745, which means that only 204 of the 800 initial VH instances make GPT-4V hallucinate. These results show that VH instances in existing image datasets are insufficient at testing MLLMs. For instance, given the results in Table 1, one may conclude that GPT-4V does not hallucinate for the shape VH mode since its accuracy is 1.000 for this VH mode. However, as we will discuss in the following, GPT-4V hallucinates substantially in the shape VH mode on our benchmarks.
@@ 4.3 Ablation Study
We use DALL·E-3 to generate VH images in Step III of our VHTest. We also evaluate other text-to-image generative models, including Midjourney 6 [15], Stable Diffusion XL 1.0 [17], and Stable Diffusion 2.1 [20] by using each of them to generate 30 VH instances, respectively. DALL·E-3 rewrites a prompt automatically to add more details, whereas the other three models lack this capability. We find that other text-to-image generative models cannot generate high-quality images when directly using the text description from Step II as a prompt. Therefore, we use GPT-4 to generate prompts given the text description for other text-to-image generative models. Specifically, we use the prompt in Figure 12 for GPT-4 to generate 30 prompts. Then, we use each text-to-image generative model to generate 30 VH images with the generated prompts. Finally, we manually craft the questions and reference answers to form 30 VH instances. Table 4 shows the accuracy of the three MLLMs on the 30 VH instances generated by different text-to-image generative models. We observe that these MLLMs achieve the lowest average accuracy of 0.144 when using DALL·E-3 in Step III. This indicates that DALL·E-3 is the most effective tool in generating VH instances that are likely to trigger VHs in MLLMs.
@@ 4.4 Mitigating VH in MLLMs
We follow LLaVA-1.5’s limited task-specific fine-tuning setting. However, we unfreeze the vision encoder since VH may result from its incorrect embedding vectors for images. We set the learning rate to 4e-6 during fine-tuning; and we fine-tune on the 960 training VH instances for one epoch (more hyperparameters and fine-tuning details are shown in Appendix E). Our fine-tuning took only 18 minutes on a single A6000.
@@
Table 6 shows the accuracy of LLaVA-1.5 and fine-tuned LLaVA-1.5 on the testing VH instances of our OEQ and YNQ benchmarks. Table 6 also shows the results on MME [3] and POPE [10], two popular existing benchmarks (not necessarily VH) to evaluate the performance of an MLLM. We find that fine-tuned LLaVA-1.5 achieves higher average accuracy than LLaVA-1.5 in both our benchmarks, while they achieve comparable results on MME and POPE benchmarks. Our results indicate that fine-tuning an MLLM on our benchmark makes it less likely to hallucinate.
@@ 5 Related Work
Hallucinations are well-known issues for generative AI, including LLMs [8, 5], MLLMs [13, 19, 25], and text-to-image generative models [24]. In general, hallucination refers to a generative model imagines factually incorrect details in its response for a given input. VH occurs when an MLLM imagines incorrect details about an image in visual question answering.
@@
Prior works have tried to benchmark MLLMs’ VHs [10, 12, 3, 25]. However, they collect VH images only from existing image datasets. This limits the diversity of VH images. Moreover, existing image datasets may have been used to pre-train an MLLM, leading to data contamination [7, 21]. Our VHTest can generate a diverse set of new VH images that do not appear in existing benchmarks. Moreover, the shape and size VH modes are formulated by us for the first time.
@@
Existing works on mitigating VHs in MLLMs can be categorized into fine-tuning-phase and testing-phase mitigation. Fine-tuning-phase mitigation focuses on improving the fine-tuning data quality [26, 12] and/or model structure [25]. These works typically freeze the vision encoder during fine-tuning, following the standard fine-tuning setting of LLaVA-1.5. We find that fine-tuning the vision encoder together reduces VHs in MLLMs. Testing-phase mitigation leverages prompt engineering with more visual evidence [9] or correction tools for hallucinated responses [29]. Testing-phase mitigation is complementary to fine-tuning-phase mitigation.
6 Conclusion and Future Work
We propose VHTest to generate VH instances to test MLLMs. We collect VH benchmarks using VHTest and we find that state-of-the-art MLLMs exhibit high hallucination rates on our benchmarks. Moreover, fine-tuning MLLMs on our benchmark reduces hallucination without sacrificing their other performance/capability. We acknowledge that our VHTest still requires human workers to manually generate a question-answer pair for an automatically generated VH image. An interesting future work is to make VHTest fully automatic so it can generate as many VH instances as needed.

### v2

@@ Abstract
[…]
tcb@breakable
Visual Hallucinations of Multi-modal Large Language Models
Wen Huangthanks: Equal contributions. University of Science & Technology of China hw2000@mail.ustc.edu.cn                        Hongbin Liu11footnotemark: 1 Duke University hongbin.liu@duke.edu
Minxin Guo The University of Hong Kong xc666@connect.hku.hk                        Neil Zhenqiang Gong Duke University neil.gong@duke.edu
@@ 1 Introduction
A multi-modal LLM (MLLM) Yang et al. (2023); Zhu et al. (2023); Chen et al. (2022); Huang et al. (2023b); Tiong et al. (2022) generates a text response for a given image and question. An MLLM typically comprises three components: a vision encoder, a vision-language connector, and an LLM. The vision encoder (e.g., CLIP Radford et al. (2021)) converts an image into an embedding vector. The vision-language connector projects an image embedding vector into the LLM’s word embedding space. The projected vector is concatenated with the token embeddings of the question to form an input to the LLM, which generates a text response.
However, MLLMs often generate text responses containing factually incorrect details about an image, known as visual hallucination (VH) Li et al. (2023); Liu et al. (2024b). Figure 1 shows an example where the MLLM hallucinates two lamps, contradicting the three lamps in the image. VHs in MLLMs pose obstacles to developing safe and trustworthy AI, which is emphasized in a recent U.S. Executive Order calling for rigorous testing to address potential harms from advanced AI systems The White House (2023).
@@
Prior works have tried to benchmark MLLMs’ VHs related to object existence Li et al. (2023); Liu et al. (2024a), optical character recognition (OCR), object counting, object positions comparing Fu et al. (2023), orientation, and viewpoint Tong et al. (2024) (concurrent to ours). However, they collect VH images only from existing image datasets like COCO Lin et al. (2014). This limits the diversity of VH images since they can only find a limited number of them. Moreover, existing image datasets may have been used to pre-train an MLLM, leading to data contamination Jacovi et al. (2023); Sainz et al. (2023). As a result, such VH images lead to a biased understanding of an MLLM’s performance, e.g., an MLLM is incorrectly concluded to perform well under VH.
@@
We propose VHTest, a tool that generates VH instances through adversarial construction to expose visual hallucinations within MLLMs. A VH instance is a triple (an image, a question, a reference answer). Our VHTest has three key steps. Step I finds initial VH instances using existing image datasets like COCO. Specifically, we first identify image pairs with high CLIP embedding similarity but low DINO v2 Oquab et al. (2023) embedding similarity. Such image pairs have contradictory similarities from two powerful vision encoders, indicating potential VHs. We note that Step I is also used in a concurrent work MMVP Tong et al. (2024). Unlike MMVP, which directly collects initial VH instances to build their benchmarks, we treat them as the raw materials for further adversarial construction. We then manually design questions and reference answers for these images to obtain initial VH instances. Step II generates a text description for a VH mode derived from the initial VH instances. A text description describes visual properties of VH images that are likely to cause MLLMs to hallucinate. Finally, Step III uses a text-to-image generative model (e.g., DALL-E 3) to generate new images based on the text descriptions. Moreover, based on some templates, we design questions and reference answers for the generated images to construct VH instances.
Using our VHTest, we construct a new benchmark dataset for evaluating VHs in MLLMs. Our benchmark contains 1,200 VH instances covering 8 VH modes. The 8 VH modes are related to existence, shape, color, orientation, OCR, size, position, and counting of visual objects in an image. Note that shape and size VH modes are formulated by us, while the other 6 VH modes were also considered in prior Yang et al. (2023) and concurrent Tong et al. (2024) studies. Figure 2 and Figure 7 show some VH instances generated by VHTest.
@@ 3.1 Step I: Finding Initial VH Instances
Since the CLIP vision encoder is the backbone of many popular MLLMs such as LLaVA Liu et al. (2023), LLaMA-Adapter Gao et al. (2023), and mPLUG-Owl Ye et al. (2023), we leverage CLIP to find initial VH instances. Our goal is to find images in an existing image dataset (e.g., COCO) that are incorrectly embedded by the CLIP vision encoder to have high similarity, despite differences in visual semantics. Such images may lead to VH for MLLMs due to their incorrect embeddings.
Specifically, given an image pair $(x_{1},x_{2})$, we compute the cosine similarity between their CLIP embedding vectors, i.e., $cos(f_{C}(x_{1}),f_{C}(x_{2}))$, where $f_{C}(\cdot)$ is an embedding vector output by CLIP. Moreover, we also use DINO v2 Oquab et al. (2023) as a reference vision encoder to compute their embedding vectors and compute their cosine similarity, i.e., $cos(f_{D}(x_{1}),f_{D}(x_{2}))$, where $f_{D}(\cdot)$ is an embedding vector output by DINO v2. Then, we find image pairs that have large cosine similarity under CLIP but small cosine similarity under DINO v2. In particular, we find image pairs that satisfy $cos(f_{C}(x_{1}),f_{C}(x_{2}))\geq 0.9$ and $cos(f_{D}(x_{1}),f_{D}(x_{2}))\leq 0.55$. Such image pairs are our candidates.
@@ 3.2 Step II: Text Description Generation
Summary version PROVIDE: {image, question, hallucinated response, detected hallucination in the response} PROMPT:   [Through the example, explain the potential causes of such hallucination and how to produce more images and questions.]
@@
Summary version PROVIDE: {image, question, hypothetical hallucinated response, detected hallucination in the hypothetical response} PROMPT:   [Through the example, explain the potential causes of the hallucination and how to generate more images and questions.]
@@ 3.3 Step III: Generating More VH Instances
Given a VH image, we prepare a question based on it. Specifically, we leverage object-driven templates to create diverse and relevant VH questions. For instance, “Describe the shape of the [object] in the picture.” is a template for the shape VH mode. We curate question templates for each VH mode and they are shown in Appendix D. Given a template for a VH mode, a human worker generates a question via manually analyzing the objects in the VH image. For instance, the human worker may replace the [object] as “pear” in the template above when the VH image contains a pear. The human worker verifies that the question should have a non-ambiguous answer based on the VH image. If no questions with non-ambiguous answers can be constructed, the VH image is discarded.
@@ 3.4 Benchmark Construction
We use VHTest to build two benchmarks. Specifically, we find initial VH instances in COCO Lin et al. (2014) and we generate 150 VH instances for each VH mode, which results in 1,200 VH instances across 8 VH modes in total. This benchmark consists of “open-ended question” (OEQ), which requires manually labeling the responses of MLLMs when testing them on the benchmark. Therefore, to facilitate automatic evaluation when testing MLLMs, we also construct a closed-ended “yes/no question” (YNQ) version of the benchmark.
Specifically, for each VH instance, we convert the open-ended question into a binary “yes/no” question. For instance, the open-ended question “Describe the shape of the pear in the picture.” is converted into “Is the shape of the pear in the picture a square?”. Moreover, the reference answer is converted into yes or no. We construct the binary questions to ensure that the YNQ benchmark has a 50/50 split of “yes”/“no” reference answers. Moreover, to ensure quality of our benchmarks, we measure the inter-annotator agreement to evaluate the consistency among different human annotators, provided in Appendix B.1.
@@ 4 Experiments
4.1 Experimental Setup
MLLMs
We evaluate three state-of-the-art MLLMs on our benchmarks: GPT-4V with its “gpt-4-vision-preview” version, LLaVA-1.5-13b Liu et al. (2023), and MiniGPT-v2  Chen et al. (2023a).
Evaluation Metric
4.2 Testing VH in MLLMs
Initial VH Instances are Insufficient
Table 1 shows the accuracy of GPT-4V on the 100 initial VH instances from COCO for each VH mode, and the number of successful initial VH instances in each VH mode. We observe that a large fraction of the initial VH instances are not successful. In particular, the average accuracy of GPT-4V across the 8 VH modes is 0.745, which means that only 204 of the 800 initial VH instances make GPT-4V hallucinate. These results show that VH instances in existing image datasets are insufficient at testing MLLMs. For instance, given the results in Table 1, one may conclude that GPT-4V does not hallucinate for the shape VH mode since its accuracy is 1.00 for this VH mode. However, as we will discuss in the following, GPT-4V hallucinates substantially in the shape VH mode on our benchmarks.
MLLMs Hallucinate on our VHTest Benchmarks
We also evaluate Gemini-pro-vision Gemini (March 2024 version), ShareGPT4V-13b Chen et al. (2023b), InstructBLIP-13b Dai et al. (2023), and Qwen-VL-Chat-7b Bai et al. (2023) on our YNQ benchmark. The results are shown in Appendix B.2. We find that these four MLLMs also suffer from visual hallucinations on our benchmark, with their accuracies all hovering around 0.54.
@@ 4.3 Ablation Study
We use DALL·E-3 to generate VH images in Step III of our VHTest. We also evaluate other text-to-image generative models, including Midjourney 6 Midjourney , Stable Diffusion XL 1.0 Podell et al. (2024), and Stable Diffusion 2.1 Rombach et al. (2022) by using each of them to generate 30 VH instances, respectively. DALL·E-3 rewrites a prompt automatically to add more details, whereas the other three models lack this capability. We find that other text-to-image generative models cannot generate high-quality images when directly using the text description from Step II as a prompt. Therefore, we use GPT-4 to generate prompts given the text description for other text-to-image generative models. Specifically, we use the prompt in Figure 13 for GPT-4 to generate 30 prompts. Then, we use each text-to-image generative model to generate 30 VH images with the generated prompts. Finally, we manually craft the questions and reference answers to form 30 VH instances. Table 4 shows the accuracy of the three MLLMs on the 30 VH instances generated by different text-to-image generative models. We observe that these MLLMs achieve the lowest average accuracy of 0.144 when using DALL·E-3 in Step III. This indicates that DALL·E-3 is the most effective tool in generating VH instances that are likely to trigger VHs in MLLMs.
@@ 4.4 Mitigating VH in MLLMs
We follow LLaVA-1.5’s limited task-specific fine-tuning setting. However, we unfreeze the vision encoder since VH may result from its incorrect embedding vectors for images. We set the learning rate to 4e-6 during fine-tuning; and we fine-tune on the 960 training VH instances for one epoch (more hyperparameters and fine-tuning details are shown in Appendix F). Our fine-tuning took only 18 minutes on a single A6000.
@@
Table 6 shows the accuracy of LLaVA-1.5 and fine-tuned LLaVA-1.5 on the testing VH instances of our OEQ and YNQ benchmarks. Table 6 also shows the results on MME Fu et al. (2023) and POPE Li et al. (2023), two popular existing benchmarks (not necessarily VH) to evaluate the performance of an MLLM. MME evaluates the perception and cognition abilities of MLLMs. The scores for MME Perception and MME Cognition are the sum of scores across the corresponding subtasks, and the total scores are 2,000 for MME Perception and 800 for MME Cognition. The POPE score represents the average F1score on random, popular, and adversarial splits of POPE. We find that fine-tuned LLaVA-1.5 achieves higher average accuracy than LLaVA-1.5 in both our benchmarks, while they achieve comparable results on MME and POPE benchmarks. Our results indicate that fine-tuning an MLLM on our benchmark makes it less likely to hallucinate. We show several qualitative comparsion results in Appendix B.3, demonstrating how the fine-tuning mitigates visual hallucinations.
@@ 5 Related Work
Hallucinations are well-known issues for generative AI, including LLMs Ji et al. (2023); Huang et al. (2023a), MLLMs Liu et al. (2024b); Rawte et al. (2023); Tong et al. (2024), and text-to-image generative model  Tong et al. (2023). In general, hallucination refers to a generative model imagines factually incorrect details in its response for a given input. VH occurs when an MLLM imagines incorrect details about an image in visual question answering.
@@
Prior works have tried to benchmark MLLMs’ VHs Li et al. (2023); Liu et al. (2024a); Fu et al. (2023); Tong et al. (2024). However, they collect VH images only from existing image datasets. This limits the diversity of VH images. Moreover, existing image datasets may have been used to pre-train an MLLM, leading to data contamination Jacovi et al. (2023); Sainz et al. (2023). Our VHTest can generate a diverse set of new VH images that do not appear in existing benchmarks. Moreover, the shape and size VH modes are formulated by us for the first time.
@@
Existing works on mitigating VHs in MLLMs can be categorized into fine-tuning-phase and testing-phase mitigation. Fine-tuning-phase mitigation focuses on improving the fine-tuning data quality Wang et al. (2024); Liu et al. (2024a) and/or model structure Tong et al. (2024). These works typically freeze the vision encoder during fine-tuning, following the standard fine-tuning setting of LLaVA-1.5. We find that fine-tuning the vision encoder together reduces VHs in MLLMs. Testing-phase mitigation leverages prompt engineering with more visual evidence Li et al. (2024) or correction tools for hallucinated responses Yin et al. (2023). Testing-phase mitigation is complementary to fine-tuning-phase mitigation.
6 Discussion
As shown in Table 2, Table 3, and Table 7, GPT-4V, LLaVA-1.5, and MiniGPT-v2 all exhibit extremely low accuracy, with an average of 0.229 on our OEQ benchmark. Moreover, all seven mainstream MLLMs achieve accuracy levels close to random guessing on our YNQ benchmark, with even the state-of-the-art GPT-4V only reaching an accuracy of 0.611. Our benchmarks effectively reveal visual hallucinations within MLLMs and are more challenging than most existing benchmarks for MLLMs.
We attribute such high challenge level to the adversarial construction idea during VH instance generation, which is similar in spirit to adversarial examples Szegedy et al. (2013) commonly used to test and improve model robustness. Unlike most existing MLLM benchmarks that are limited to the image space of existing image datasets like COCO, VHTest derives text descriptions from initial VH instances, which play a role of “adversarial direction”. Then VHTest exploits text-to-image generative models to generate challenging VH instances following such adversarial direction. This moves beyond the image space of limited existing image datasets and into the manifold of text-to-image generative models, thereby making it possible to generate more challenging VH instances. Thus, our VHTest offers a new perspective for constructing MLLM benchmarks for future research.
7 Conclusion
We propose VHTest to generate VH instances to test MLLMs. We collect VH benchmarks using VHTest and we find that state-of-the-art MLLMs exhibit high hallucination rates on our benchmarks. Moreover, fine-tuning MLLMs on our benchmark reduces hallucination without sacrificing their other performance/capability.
8 Limitations and Future Work
We acknowledge that our VHTest still requires human workers to manually generate a question-answer pair for an automatically generated VH image. An interesting future work is to make VHTest fully automatic so it can generate as many VH instances as needed.